OpenStudy (anonymous):

Find the slope of the line parallel to the line y = -5x + 24 and goes through the point (24,79).

OpenStudy (anonymous):

Do you know what point-slope form is?

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

Point-slope form is \[y-y_1=m(x-x_1)\] . Sooooooo.... you need to put the two points in this equation. So \[y-79=-5(x-24)\]

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

Then you need to simplify that and put it into slope-intercept form which is y=mx+b

OpenStudy (anonymous):

okay thank you

OpenStudy (anonymous):

would it be y=-5x+199

OpenStudy (anonymous):

Yep